HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ACTION SPECTROSCOPY STUDY OF (H_{2}O-X)^{+} COMPLEXES IN THE OVERTONE RANGE

\underline{J. FRÉREUX} , B. TOMASETTI , G. VANLANCKER , A. DELCORTE , X. URBAIN and C. LAUZIN, Institute of Condensed Matter and Nanosciences,

Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

H_{2}O^{+} plays a role in the ion chemistry of the lower atmosphere and acts as a nucleation center^{1}. In this talk I will present new results concerning (H_{2}O-X)^{+} complexes, with X = Ar, CO_{2} and N_{2}O. These complexes are produced using a supersonic expansion and an electron beam. Photo-fragmentation spectroscopy is performed with an OPO laser in the near-IR (1000-2100 nm) and by monitoring the signal of H_{2}O^{+} fragments as a function of the laser wavelength. New vibrational bands will be reported. The results concerning (H_{2}O-Ar)^{+} will be compared with previous results obtained in the literature^{2}. Tentative assignment of the observed bands will be performed by a careful comparison between the different complexes. Finally, I will present the status of our efforts in the production of very large ionic clusters, with pure (Ar)_{n}^{+} and mixed clusters (H_{2}O_{m}-Ar_{n})^{+} containing up to thousands of units.

1doi:10.1021/cr5003479, Shuman, N. S., Hunton, D. E., & Viggiano, A. A. (2015). Ambient and modified atmospheric ion chemistry: From top to bottom. Chemical reviews, 115(10), 4542-4570.

2doi:10.1063/1.4998419, Wagner, J. P., McDonald, D. C., & Duncan, M. A. (2017). Near-infrared spectroscopy and anharmonic theory of the H2O+ Ar1, 2 cation complexes. The Journal of chemical physics, 147(10), 104302.

p-number: p182  Submitted on Thu, 15 Jun 2023 14:29:29 +0200